# A Review of LLM-Assisted Ideation

### v2

@@ 4 The Hourglass Ideation Framework for LLM-assisted Ideation
To understand the roles and contributions of large language models (LLMs) in supporting ideation, we propose the Hourglass Ideation Framework (Figure 3), an iterative model that captures the key stages and interactions involved in leveraging LLMs for creative ideation. The framework draws inspiration from the structuring of ideation factors proposed by Gerlach and Brem [gerlach2017idea], but is derived from our comprehensive survey of LLM-assisted ideation literature. It adapts and extends the existing model to align with the unique characteristics and requirements of LLM-based ideation processes.
Figure 3: The Hourglass Ideation Framework for LLM-assisted Ideation. The hourglass shape of the framework visualizes the divergent and convergent dynamics of the ideation process, which consists of three overarching phases: preparation (blue), divergent (orange), and convergent (red). Within these phases, there are seven key stages: scope definition, foundational materials collection, foundational materials structuring, idea generation, idea refinement, idea evaluation, and multi-idea evaluation and selection. The framework emphasizes the iterative nature of the ideation process, with arrows indicating potential loops between stages. The bars show the involvement of LLMs and non-LLMs methods in each stage, while the percentages represent the proportion of reviewed papers that included each method.
The Hourglass Ideation Framework (Figure 3) comprises seven key stages: scope definition, foundational materials collection, foundational materials structuring, idea generation, idea refinement, idea evaluation, and multi-idea evaluation and selection. These stages are organized into three overarching phases: preparation, divergent, and convergent. The preparation phase establishes the groundwork for ideation by defining objectives and curating relevant resources. The divergent phase focuses on the generation and exploration of a wide range of ideas, while the convergent phase narrows down and refines the most promising concepts.
@@ 4.1 An Iterative Framework for LLM Assistance
    Foundational Materials Structuring - This materials structuring stage focuses on organizing and synthesizing the diverse texts and data sources gathered during the previous stage. It plays a pivotal role in the LLM-assisted ideation process, as it helps researchers systematically analyze and refine the input resources, facilitating their effective utilization in subsequent ideation stages. LLMs contribute to this process primarily through text retrieval, structured encoding, ranking, and filtering.
@@ 4.2 Overview of LLM use in Ideation Stages
[…]
@@
[…]
@@
[…]
@@ 4.3.1 Scope Definition
In 53 out of 59 studies, the ideation scope was determined by the organizers or researchers, ensuring alignment with the study’s goals and research questions. However, in 3 studies related to writing tasks [goodman2022lampost_W1, kim2024towards_W2, benharrak2024writer_W14], participants were allowed to define their own scope by providing specific writing prompts or objectives. This approach highlights the potential for user-driven scope definition in certain ideation contexts.
Notably, only 3 studies (5%) employed LLMs to assist in the scope definition process. In study [pu2024ideasynth_S1], the LLM helped refine the initial research problem by generating clarifying questions or proposing alternative scopes based on relevant literature from the pre-ideation foundations. Similarly, study [liu2024personaflow_S2] utilized an LLM to modify the user’s initial research question into a revised problem statement. In study [zhang2023visar_W4], the LLM provided hierarchical writing goal suggestions to guide the scope definition for argumentative writing tasks.
@@ 4.3.2 Foundational Materials Collection
1) Background information, such as existing solutions [junior2023chatgpt_D11, veloso2024forming_D12], market conditions [bilgram2023accelerating_D2], competitor information [xu2024jamplate_G3], and scenario-based context [paredes2024creative_G17];
[…]
3) Data collections, such as biology databases [chen2024asknaturenet_G10, kang2024biospark_G14], patent repositories [lee2024generating_D7], specialized document corpora [yuan2024llmcrit_W15], engineering prototype datasets [ege2024chatgpt_D10], custom code libraries [qian2024shape_M5], and comprehensive research resources [blazevic2024real_G15, chen2024use_S5];
4) Guidelines and criteria, including ecological design principles [lee2024generating_D7], authoritative references [yuan2024llmcrit_W15], and standards for generating creative feedback [yuan2024llmcrit_W15];
5) Personas, representing diverse perspectives such as product testers [schuller2024generating_G9], designers, users, and products [kocaballi2023conversational_D1], as well as customized roles for providing writing feedback [benharrak2024writer_W14] and evaluating scientific ideas [liu2024personaflow_S2].
[…]
These findings suggest that LLMs can play a valuable role in curating, generating, and processing pre-ideation content, helping to establish a rich and diverse foundation for creative exploration. By leveraging their ability to analyze and summarize large volumes of data, LLMs can surface relevant information, identify patterns and connections, and distill complex materials into actionable insights.
@@ 4.3.3 Foundational Materials Structuring
Foundational Materials xxx Structuring x Non-LLMs Similarity [chen2024asknaturenet_G10], [kang2024biospark_G14], [blazevic2024real_G15], [lee2024generating_D7], [wang2024lave_M4], [qian2024shape_M5], [liu2024personaflow_S2]
Clustering [lee2024generating_D7], [kang2024biospark_G14]
Ranking [blazevic2024real_G15], [lee2024generating_D7]
Filtering [lee2024generating_D7]
Sorting [kang2024biospark_G14]
LLMs Retrieval [pu2024ideasynth_S1], [radensky2024scideator_S3], [blazevic2024real_G15], [qian2024shape_M5], [lee2024generating_D7]
Structuring [chen2024asknaturenet_G10], [kang2024biospark_G14]
Ranking [wang2024lave_M4]
[…]
@@
1) Similarity calculation: This method involves measuring the semantic relatedness between textual elements, such as documents or concepts, to identify patterns and connections within the collected materials. Similarity metrics, such as cosine similarity [chen2024asknaturenet_G10, lee2024generating_D7, wang2024lave_M4, liu2024personaflow_S2] and Euclidean distance [lee2024generating_D7], were used to compute the similarity between embeddings of textual data.
2) Clustering: Clustering methods were used to group similar textual elements together based on their semantic or structural properties. Study [lee2024generating_D7] applied text clustering to group semantically similar guidelines generated by LLMs, making them more accessible and user-friendly.
3) Ranking: Non-LLM ranking techniques were applied to prioritize and order the collected materials based on their relevance or importance to the ideation context. For example, the study [blazevic2024real_G15] employed a combination of topic modeling and similarity algorithms to rank scientific publications.
4) Filtering: Filtering methods were used to refine the collected materials by removing irrelevant or redundant information. Study [lee2024generating_D7] applied regular expressions and predefined lists to filter out inaccurate principles generated by LLMs.
5) Sorting: Sorting techniques were employed to arrange the collected materials in a specific order based on predefined criteria. For instance, study [kang2024biospark_G14] sorted taxonomic ranks in an increasing order based on the number of immediate children nodes to identify sparsely populated branches for expansion.
@@
1) Text retrieval based on semantic similarity: Several studies employed LLMs to identify and retrieve relevant text segments from large corpora by calculating semantic similarity between the query and the stored embeddings [pu2024ideasynth_S1, radensky2024scideator_S3, blazevic2024real_G15, qian2024shape_M5, lee2024generating_D7]. This approach, known as Retrieval-Augmented Generation (RAG), is a cutting-edge technique that combines the strengths of LLMs and information retrieval systems to generate more accurate and relevant outputs.
2) Structured encoding of text data: LLMs were utilized to encode textual data into structured formats, such as knowledge graphs or taxonomic hierarchies [chen2024asknaturenet_G10, kang2024biospark_G14]. This structured representation facilitates the exploration and analysis of complex relationships within the data, enhancing the ideation process.
3) Ranking: LLMs were employed to rank the relevance of video narrations based on their semantic similarity to a given query [wang2024lave_M4]. This ranking mechanism enables users to quickly identify the most pertinent content for their ideation needs.
[…]
However, our review identified a lack of research exploring the use of LLMs for similarity calculation and clustering in the stage of foundational materials structuring. These techniques have been widely used in non-LLM methods, as demonstrated by several studies [chen2024asknaturenet_G10, kang2024biospark_G14, lee2024generating_D7, liu2024personaflow_S2]. The absence of LLM-based similarity and clustering approaches presents an opportunity for future research to investigate how LLMs can be leveraged to enhance the effectiveness and efficiency of these structuring techniques. By incorporating LLMs’ semantic understanding capabilities, researchers may be able to develop more sophisticated and context-aware methods for grouping and organizing foundational materials.
@@ 4.3.5 Idea Refinement
Idea xxxxxx Refinement Non-LLMs Individual 111 1 idea refinement Modification [xu2024jamplate_G3], [huang2023causalmapper_G5], [hou2024c2ideas_D13], [huang2024plantography_D14], [yuan2024llmcrit_W15], [pu2024ideasynth_S1], [liu2024personaflow_S2]
Elimination [huang2023causalmapper_G5], [radensky2024scideator_S3]
Combination Combination [shaer2024ai_G1], [suh2024luminate_G13]
LLMs Individual 111 1 idea refinement Expansion [di2022idea_G4], [aikawa2023introducing_G8], [suh2024luminate_G13], [kocaballi2023conversational_D1], [yuan2022wordcraft_W5], [ghajargar2022redhead_W7], [pu2024ideasynth_S1], [cai2023designaid_M1], [brade2023promptify_M3], [wang2024lave_M4]
Modification [di2022idea_G4], [bhavya2023cam_G16], [lanzi2023chatgpt_D6], [ege2024chatgpt_D10], [junior2023chatgpt_D11], [goodman2022lampost_W1], [reza2024abscribe_W13], [benharrak2024writer_W14], [lee2024gpt_S6], [qian2024shape_M5]
Substitution [lanzi2023chatgpt_D6], [ege2024chatgpt_D10], [zhang2023visar_W4], [yuan2022wordcraft_W5], [reza2024abscribe_W13], [pu2024ideasynth_S1]
Rewriting [di2022idea_G4], [gonzalez2024collaborative_G7], [goodman2022lampost_W1], [yuan2022wordcraft_W5], [qian2024shape_M5], [chen2024foundation_M8]
Elimination [blazevic2024real_G15], [chung2022talebrush_W3]
Multiple perspectives [xu2024jamplate_G3], [aikawa2023introducing_G8], [suh2024luminate_G13], [kang2024biospark_G14], [zhang2023visar_W4], [benharrak2024writer_W14], [pu2024ideasynth_S1], [liu2024personaflow_S2], [radensky2024scideator_S3], [wang2023popblends_M2]
Questioning [xu2024jamplate_G3], [bilgram2023accelerating_D2], [wang2023task_D8], [kim2024towards_W2]
Combination Combination [di2022idea_G4], [besta2024graph_G12], [kang2024biospark_G14], [lanzi2023chatgpt_D6], [petridis2023anglekindling_W6], [chen2024foundation_M8]
@@ 4.3.6 Idea Evaluation
LLM-based evaluation methods primarily include providing feedback or suggestions on product prototypes [ege2024chatgpt_D10, junior2023chatgpt_D11, chen2024foundation_M8] or text [yuan2024llmcrit_W15, kim2024towards_W2, zhang2023visar_W4, goldi2024intelligent_W10, pu2024ideasynth_S1]; generating virtual personas to offer feedback on academic research ideas [liu2024personaflow_S2], written content [benharrak2024writer_W14], or engaging in virtual interviews to provide feedback [kocaballi2023conversational_D1]; integrating LLMs into workflows to facilitate task evaluation [dharuman2023protein_D5, lee2024gpt_S6]; scoring [bhavya2023cam_G16, besta2024graph_G12] and rating ideas [shaer2024ai_G1, aikawa2023introducing_G8] based on specific criteria or functions. Additionally, one study [radensky2024scideator_S3] employed similarity-based reordering of idea-related reference materials to assess the novelty of ideas.
Several studies evaluated ideas using general assessment tools such as strengths and weaknesses [rosenberg2023conversational_G6, kang2024biospark_G14] or specific evaluation criteria, including relevance, innovation, and insightfulness [shaer2024ai_G1]; unique value, advantages, and disadvantages [xu2024jamplate_G3]; originality, efficacy, and feasibility [aikawa2023introducing_G8]; and analogical style, meaningfulness, and novelty [bhavya2023cam_G16].
@@ 4.3.7 Multi-idea Evaluation and Selection
Multi-idea xxx Evaluation x x and Selection Non-LLMs Comparison [shaer2024ai_G1], [huang2023causalmapper_G5], [suh2024luminate_G13], [lanzi2023chatgpt_D6], [hou2024c2ideas_D13], [goodman2022lampost_W1], [chung2022talebrush_W3], [zhang2023visar_W4], [yuan2022wordcraft_W5], [kim2023cells_W9], [goldi2024intelligent_W10], [reza2024abscribe_W13], [benharrak2024writer_W14], [pu2024ideasynth_S1], [liu2024personaflow_S2], [radensky2024scideator_S3], [cai2023designaid_M1], [brade2023promptify_M3], [chen2024foundation_M8]
Filtering [di2022idea_G4], [huang2023causalmapper_G5], [suh2024luminate_G13], [dharuman2023protein_D5], [radensky2024scideator_S3], [chen2024foundation_M8]
Similarity [bhavya2023cam_G16], [kim2023cells_W9], [cai2023designaid_M1], [wang2023popblends_M2]
Clustering [gonzalez2024collaborative_G7], [suh2024luminate_G13], [brade2023promptify_M3], [kim2023cells_W9]
Sorting [besta2024graph_G12], [suh2024luminate_G13]
Ranking [besta2024graph_G12], [wang2023popblends_M2]
Searching [suh2024luminate_G13]
LLMs Comparison [xu2024jamplate_G3], [kang2024biospark_G14]
Ranking [blazevic2024real_G15], [bhavya2023cam_G16]
Rating [bhavya2023cam_G16]
LLMs contribute to the multi-idea evaluation and selection process through three primary methods: comparing multiple ideas [xu2024jamplate_G3, kang2024biospark_G14], ranking multiple ideas [blazevic2024real_G15, bhavya2023cam_G16], and rating multiple ideas [bhavya2023cam_G16] to facilitate idea selection.

### v3

@@ 4 The Hourglass Ideation Framework for LLM-assisted Ideation
To understand the roles and contributions of large language models (LLMs) in supporting ideation, we propose the Hourglass Ideation Framework (Figure 3), an iterative model that captures the key stages and interactions involved in leveraging LLMs for creative ideation. The framework draws inspiration from the structuring of ideation factors proposed by Gerlach [32], but is derived from our comprehensive survey of LLM-assisted ideation literature. It adapts and extends the existing model to align with the unique characteristics and requirements of LLM-based ideation processes.
Figure 3: The Hourglass Ideation Framework for LLM-assisted Ideation. The hourglass shape of the framework visualizes the divergent and convergent dynamics of the ideation process, which consists of three overarching phases: preparation (blue), divergent (orange), and convergent (red). Within these phases, there are seven key stages: scope definition, foundational materials collection, foundational materials structuring, idea generation, idea refinement, idea evaluation, and multi-idea evaluation and selection. The framework emphasizes the iterative nature of the ideation process, with arrows indicating potential loops between stages. The bars show the involvement of LLMs and non-LLMs methods in each stage. The percentages and numbers represent the proportion and quantity of reviewed papers, respectively.
The Hourglass Ideation Framework (Figure 3) comprises seven key stages which are organized into three overarching phases: preparation, divergent, and convergent. The preparation phase, represented by the narrow top of the hourglass, establishes the groundwork for ideation by defining objectives and curating relevant resources. The divergent phase, forming the expanding middle section, focuses on generating and exploring a wide range of ideas, while the convergent phase, mirroring the narrow bottom, narrows down and refines the most promising concepts.
@@ 4.1 An Iterative Framework for LLM Assistance
    Foundational Materials Structuring - This materials structuring stage focuses on organizing and synthesizing the diverse texts and data sources gathered during the previous stage. It plays a pivotal role in the LLM-assisted ideation process, as it helps researchers systematically analyse and refine the input resources, facilitating their effective utilization in subsequent ideation stages. LLMs contribute to this process primarily through text retrieval, structured encoding, ranking, and filtering.
@@ 4.2 Overview of LLM use in Ideation Stages
From the 61 papers we reviewed, Table 3 presents an overview of how LLMs have been used within different ideation stages. Note that the techniques described cover both LLM($\blacksquare$) and non-LLM($\square$) methods. Stages annotated with both symbols ($\square\blacksquare$) indicate that a stage is performed jointly using LLM and non-LLM techniques.
@@
Table 3: Large language models’ (LLMs) contributions to the critical ideation stages. Includes LLM($\blacksquare$), Non-LLM($\square$) and both($\square\blacksquare$) methods.
@@
[95] General G1 $\square\phantom{\blacksquare}$ - - $\square\blacksquare$□ ■ $\square\phantom{\blacksquare}$ $\square\blacksquare$□ ■ $\square\phantom{\blacksquare}$ -
[…]
[87] General G13 $\square\phantom{\blacksquare}$ - - $\square\blacksquare$□ ■ $\square\blacksquare$□ ■ $\square\phantom{\blacksquare}$ $\square\phantom{\blacksquare}$ 7 -→ 4 / 5
[…]
[92] General G17 - $\square\phantom{\blacksquare}$ - $\square\blacksquare$□ ■ - - - -
@@
[48] Writing W8 $\square\phantom{\blacksquare}$ - - $\square\blacksquare$□ ■ - - - -
[…]
[55] Writing W10 $\square\phantom{\blacksquare}$ - - $\square\blacksquare$□ ■ - $\square\blacksquare$□ ■ $\square\phantom{\blacksquare}$ -
@@
[70] Science S4 $\square\phantom{\blacksquare}$ - - $\square\blacksquare$□ ■ - - - -
@@ 4.3.1 Scope Definition
In 53 out of 59 studies, the ideation scope was determined by the organizers or researchers, ensuring alignment with the study’s goals and research questions. However, in 3 studies related to writing tasks [56, 43, 46], participants were allowed to define their own scope by providing specific writing prompts or objectives. This approach highlights the potential for user-driven scope definition in certain ideation contexts.
Notably, only 3 studies (5%) employed LLMs to assist in the scope definition process. In study [67], the LLM helped refine the initial research problem by generating clarifying questions or proposing alternative scopes based on relevant literature from the pre-ideation foundations. Similarly, study [68] utilized an LLM to modify the user’s initial research question into a revised problem statement. In study [54], the LLM provided hierarchical writing goal suggestions to guide the scope definition for argumentative writing tasks.
@@ 4.3.2 Foundational Materials Collection
1) Background information, such as existing solutions [73, 75], market conditions [36], competitor information [83], and scenario-based context [92];
2) Ideation materials, including textual prompts for writing inspiration [56, 43], video clips [66], audio recordings of group discussions [94], story elements for picture book creation [63], and scholarly literature for exploring research ideas [67, 68, 69];
3) Data collections, such as biology databases [86, 61], patent repositories [41], specialized document corpora [81], engineering prototype datasets [37], custom code libraries [64], and comprehensive research resources [90, 71];
4) Guidelines and criteria, including ecological design principles [41], authoritative references [81], and standards for generating creative feedback [81];
5) Personas, representing diverse perspectives such as product testers [89], designers, users, and products [35], as well as customized roles for providing writing feedback [46] and evaluating scientific ideas [68].
Among the studies that incorporated the foundational materials collection stage, 16 (67%) utilized LLMs to help generate or summarize foundational content. They contribute to this stage in two ways. First, they can generate background [83, 36, 75], guidelines [41], criteria [81], and personas [89, 35, 46, 68] that guide and inspire the ideation process. Second, LLMs can summarize and synthesize ideation materials [94, 56, 43, 66, 63] and scholarly literature [67, 68, 69], extracting key insights to inform the creative process.
These findings suggest that LLMs can play a valuable role in curating, generating, and processing pre-ideation content, helping to establish a rich and diverse foundation for creative exploration. By leveraging their ability to analyse and summarize large volumes of data, LLMs can surface relevant information, identify patterns and connections, and distill complex materials into actionable insights.
@@ 4.3.3 Foundational Materials Structuring
Foundational Materials xxx Structuring x Non-LLMs Similarity [86], [61], [90], [41], [66], [64], [68]
Clustering [41], [61]
Ranking [90], [41]
Filtering [41]
Sorting [61]
LLMs Retrieval [67], [69], [90], [64], [41]
Structuring [86], [61]
Ranking [66]
Filtering [63]
@@
1) Similarity calculation: This method involves measuring the semantic relatedness between textual elements, such as documents or concepts, to identify patterns and connections within the collected materials. Similarity metrics, such as cosine similarity [86, 41, 66, 68] and Euclidean distance [41], were used to compute the similarity between embeddings of textual data.
2) Clustering: Clustering methods were used to group similar textual elements together based on their semantic or structural properties. Study [41] applied text clustering to group semantically similar guidelines generated by LLMs, making them more accessible and user-friendly.
3) Ranking: Non-LLM ranking techniques were applied to prioritize and order the collected materials based on their relevance or importance to the ideation context. For example, the study [90] employed a combination of topic modeling and similarity algorithms to rank scientific publications.
4) Filtering: Filtering methods were used to refine the collected materials by removing irrelevant or redundant information. Study [41] applied regular expressions and predefined lists to filter out inaccurate principles generated by LLMs.
5) Sorting: Sorting techniques were employed to arrange the collected materials in a specific order based on predefined criteria. For instance, study [61] sorted taxonomic ranks in an increasing order based on the number of immediate children nodes to identify sparsely populated branches for expansion.
@@
1) Text retrieval based on semantic similarity: Several studies employed LLMs to identify and retrieve relevant text segments from large corpora by calculating semantic similarity between the query and the stored embeddings [67, 69, 90, 64, 41]. This approach, known as Retrieval-Augmented Generation (RAG), is a cutting-edge technique that combines the strengths of LLMs and information retrieval systems to generate more accurate and relevant outputs.
2) Structured encoding of text data: LLMs were utilized to encode textual data into structured formats, such as knowledge graphs or taxonomic hierarchies [86, 61]. This structured representation facilitates the exploration and analysis of complex relationships within the data, enhancing the ideation process.
3) Ranking: LLMs were employed to rank the relevance of video narrations based on their semantic similarity to a given query [66]. This ranking mechanism enables users to quickly identify the most pertinent content for their ideation needs.
4) Filtering data information: LLMs were used to filter and refine textual data, such as extracting relevant keywords and descriptions from story scripts [63]. This filtering process helps to distill the most valuable information from the collected materials, ensuring a more focused and effective ideation process.
However, our review identified a lack of research exploring the use of LLMs for similarity calculation and clustering in the stage of foundational materials structuring. These techniques have been widely used in non-LLM methods, as demonstrated by several studies [86, 61, 41, 68]. The absence of LLM-based similarity and clustering approaches presents an opportunity for future research to investigate how LLMs can be leveraged to enhance the effectiveness and efficiency of these structuring techniques. By incorporating LLMs’ semantic understanding capabilities, researchers may be able to develop more sophisticated and context-aware methods for grouping and organizing foundational materials.
@@ 4.3.5 Idea Refinement
Idea xxxxxx Refinement Non-LLMs Individual 111 1 idea refinement Modification [83], [74], [40], [62], [81], [67], [68]
Elimination [74], [69]
Combination Combination [95], [87]
LLMs Individual 111 1 idea refinement Expansion [88], [91], [87], [35], [47], [51], [67], [57], [59], [66]
Modification [88], [85], [80], [37], [73], [56], [45], [46], [72], [64]
Substitution [80], [37], [54], [47], [45], [67]
Rewriting [88], [93], [56], [47], [64], [60]
Elimination [90], [50]
Multiple perspectives [83], [91], [87], [61], [54], [46], [67], [68], [69], [58]
Questioning [83], [36], [78], [43]
Combination Combination [88], [82], [61], [80], [49], [60]
@@ 4.3.6 Idea Evaluation
LLM-based evaluation methods primarily include providing feedback or suggestions on product prototypes [37, 73, 60] or text [81, 43, 54, 55, 67]; generating virtual personas to offer feedback on academic research ideas [68], written content [46], or engaging in virtual interviews to provide feedback [35]; integrating LLMs into workflows to facilitate task evaluation [42, 72]; scoring [85, 82] and rating ideas [95, 91] based on specific criteria or functions. Additionally, one study [69] employed similarity-based reordering of idea-related reference materials to assess the novelty of ideas.
Several studies evaluated ideas using general assessment tools such as strengths and weaknesses [94, 61] or specific evaluation criteria, including relevance, innovation, and insightfulness [95]; unique value, advantages, and disadvantages [83]; originality, efficacy, and feasibility [91]; and analogical style, meaningfulness, and novelty [85].
@@ 4.3.7 Multi-idea Evaluation and Selection
Multi-idea xxx Evaluation x x and Selection Non-LLMs Comparison [95], [74], [87], [80], [40], [56], [50], [54], [47], [44], [55], [45], [46], [67], [68], [69], [57], [59], [60]
Filtering [88], [74], [87], [42], [69], [60]
Similarity [85], [44], [57], [58]
Clustering [93], [87], [59], [44]
Sorting [82], [87]
Ranking [82], [58]
Searching [87]
LLMs Comparison [83], [61]
Ranking [90], [85]
Rating [85]
LLMs contribute to the multi-idea evaluation and selection process through three primary methods: comparing multiple ideas [83, 61], ranking multiple ideas [90, 85], and rating multiple ideas [85] to facilitate idea selection.